OpenStudy (anonymous):

5. What is an equation of the line in slope intercept form? M=2/7 and the y intercept is (0,-12)

OpenStudy (anonymous):

http://www.mathopenref.com/coordequation.html

OpenStudy (ivettef365):

the formula is y = mx + b m= slope b = y-int you have both numbers that you need

OpenStudy (anonymous):

Ok. But i dont know how to find the answer..

OpenStudy (anonymous):

Plug in the numbers

OpenStudy (anonymous):

IF the answer was a picture of a graph then it would make sense. But here are the choices. A. Y=2/7x+12 B. y=2/7x-12 C. y=12x-2/7 D. y=12x+2/7

OpenStudy (anonymous):

IM confused

OpenStudy (anonymous):

WAIT!

OpenStudy (anonymous):

I understand. Thank yall!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

THe answer is A right?

OpenStudy (anonymous):

should be b

OpenStudy (anonymous):

Isn't the 12 a negative ? It should be B .

OpenStudy (anonymous):

OHHHH. I wasnt paying attention to the negative

OpenStudy (anonymous):

Haha , make sure to pay attention !

OpenStudy (anonymous):

It's critical if you want to get the correct answer .